OpenStudy (anonymous):

Find the derivative of the function by the limit process. f(x) = 7 x

OpenStudy (emmatassone):

If you mean f(x) =7x, i can give you a hand: To find the derivative with the definition you always have to propose this limit: \[\lim_{h \rightarrow 0} \frac{ f(x+h) - f(x) }{ h }\] In this case we know that f(x)=7x\[\lim_{h \rightarrow 0}\frac{ 7(x+h) - 7x }{ h }\] from there try to work with this limit to get the solution using the properties you know.

OpenStudy (irishboy123):

\[=\lim_{h \rightarrow 0}\frac{ 7x+7h - 7x }{ h }\]